1) Determine the stopping distances for a car with an initial speed of 97km/h and human reaction time of 1.0s , for an acceleration a=?3.8m/s2.

2) Determine the stopping distances for a car with an initial speed of 97

km/h and human reaction time of 1.0s , for an acceleration a=?8.2m/s2.
Answer #1

u = 97 km / hr = 26.944 m/s

a = -3.8 m/s^2

the distance covered in 1 sec is 26.944 m

then by using the thord equation of motion

v^2 - u^2 = 2 * a * s

0 - 26.944^2 = 2 * -3.8 * s

s = - 26.944 / 2 * -3.8

s = 95.52 m

then the total stopping distance is 95.52 + 26.944 = 122.464 m

2)

the distance covered in 1 sec is 26.944 m

then by using the third equation of motion

v^2 - u^2 = 2 * a * s

0 - 26.944^2 = 2 * -8.2 * s

s = 44.26m

so the total stopping distance covered is 44.26 + 26.944 = 71.2 m